# Sample-efficient diffusion-based control of complex nonlinear systems

### v1

@@ 5.3 Ablation Study
Overall ablation study. We explore the main performance against each ablation of the original SEDC. Specifically, w/o DSD removes the inverse dynamics, unifying the diffusion of system state and control input, i.e. $\mathbf{x}=[\mathbf{u},\mathbf{y}]$. Therefore, the diffusion model is required to simultaneously capture the temporal information and implicit dynamics of the control and system trajectory. Note that the inpainting mechanism and gradient guidance are retained. w/o DMD removes the decomposition design, resulting in a single 1-D Unet structure as the denoising network, following DecisionDiff (Ajay et al., 2022). Finally, w/o GSF reports the performance without iterative self-finetuning, which means the model only uses the original dataset to train itself. To show the sample-efficiency performance, we also investigate the results under less amount of training sample(10%). For w/o DMD and w/o DSD, we adjust the number of trainable parameters at a comparable level against the original version.
@@
Table 2: SEDC and w/o DSD Target Loss comparison across different state dimensions in Kuramoto. The Dec. indicates the reduction in target loss achieved by SEDC compared to w/o DSD. Notably, the magnitude of loss reduction increases proportionally with network dimensionality, demonstrating that DSD’s performance enhancement scales positively with dimensional growth.
$N$ 4 5 6 7 8
SEDC (e-6) 3.45 2.89 5.67 4.12 8.90
w/o DSD (e-4) 3.98 3.23 16.78 12.45 54.32
Dec.(%) 99.13 99.11 99.66 99.67 99.84
Effectiveness of DSD. To evaluate DSD’s effectiveness against the curse of dimensionality, we compared the performance of original and w/o DSD models across Kuramoto systems with dimensions ranging from $N=4$ to $N=8$. Experimental results (Table 2) show that performance degradation (Dec.) from w/o DSD increases with system dimensionality, demonstrating DSD’s enhanced effectiveness in higher-dimensional systems and validating its capability to address dimensionality challenges.
To investigate the effectiveness of dynamical learning, we compared the consistency between action sequences and diffusion-sampled state trajectories in models with and without DSD. While both approaches can sample state trajectories from diffusion samples, they differ in action generation: SEDC uses inverse dynamics prediction, whereas w/o DSD obtains actions directly from diffusion samples by simultaneously diffusing states and control inputs. We test both models using identical start-target conditions and visualize the state induced from the generated actions and the state sampled from the diffusion model, along with the difference (error) between the above two states in Figure 4. We can observe that SEDC’s action-induced state trajectories showed significantly higher consistency with sampled trajectories compared to w/o DSD, demonstrating that DSD using inverse dynamics achieves more accurate learning of control-state dynamical relationships.
Table 3: Performance degradation using different denoiser output with varying nonlinearity strength $\gamma$ in the Kuramoto system. The Dec. indicates the reduction in target loss achieved by nonlinear output $\mathbf{O}_{1}+\mathbf{O}_{2}$ compared to linear output $\mathbf{O}_{1}$. Notably, the magnitude of loss reduction increases proportionally with the nonlinearity strength $\gamma$, indicating that the quadratic term exhibits enhanced capability in capturing nonlinear dynamics as the system’s nonlinearity intensifies.
@@
Effectiveness of DMD. To investigate the contribution of DMD’s dual-Unet architecture to nonlinearity learning, we conducted experiments on the Kuramoto system with varying degrees of nonlinearity (controlled by the coefficient $\gamma\in\{1,2,4\}$ of the nonlinear sinusoidal term, where larger values indicate stronger nonlinearity). We compared the performance between using only the linear intermediate output ($\hat{\mathbf{x}}_{0}=\mathbf{O}_{1}$) of the denoising network and the original nonlinear output ($\hat{\mathbf{x}}_{0}=\mathbf{O}_{1}+\mathbf{O}_{2}$) in terms of Target Loss. As shown in Table LABEL:tab:dmd, the performance degradation (Dec.) from using only $\mathbf{O}_{1}$ becomes more pronounced as nonlinearity increases. This demonstrates both the significance of the nonlinear branch $\mathbf{O}_{2}$ in capturing strong nonlinear dynamics and the effectiveness of decoupling linear and nonlinear modes in handling system nonlinearity.
Effectiveness of GSF. To validate GSF’s effectiveness in guiding the model toward learning the optimal (energy-efficient) target distribution, we conduct ablation studies on fine-tuning rounds and evaluate control signal energy on the test set. Results show decreasing energy metrics over rounds, confirming our approach’s convergence behavior toward near-optimal control strategies. Due to space limits, results and detailed discussion are provided in Appendix F.
@@ 6 Conclusion
In this paper, we presented SEDC, a novel sample-efficient diffusion-based framework for complex nonlinear system control. By addressing fundamental challenges in data-driven control through three key innovations - Decoupled State Diffusion (DSD), Dual-Mode Decomposition (DMD), and Guided Self-finetuning (GSF) - SEDC achieves superior control performance while significantly reducing sample requirements. Our comprehensive experiments across three nonlinear systems demonstrate that SEDC outperforms existing methods by 39.5%-49.4% in control accuracy while maintaining computational efficiency. Most notably, SEDC achieves state-of-the-art performance using only 10% of the training samples required by baseline methods, marking a significant advancement in sample-efficient control of complex systems. These results validate our approach’s effectiveness in addressing the curse of dimensionality, handling strong nonlinearities, and bridging the gap between non-optimal training data and optimal control solutions. As complex system control continues to evolve across various domains, SEDC’s sample-efficient framework provides a promising direction for future research and practical applications in data-driven control.
@@ Appendix A Algorithm form of SEDC
Predict clean states: $\hat{\mathbf{y}}^{0}=\mathcal{G}_{\theta}(\mathbf{y}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})$
Predict controls: $\hat{\mathbf{u}}_{t}=f_{\phi}(\hat{\mathbf{y}}_{t}^{0},\hat{\mathbf{y}}_{t+1}^{0})$
Compute losses:  $L_{\text{diff}}=\|\mathbf{y}-\hat{\mathbf{y}}^{0}\|^{2}$
@@
             Predict $\hat{\mathbf{y}}^{0}=\mathcal{G}_{\theta}(\mathbf{y}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})$
Compute gradient: $g=\nabla_{\mathbf{y}^{k}}J(\hat{\mathbf{y}}^{0})$
@@
Enforce constraints: $\mathbf{y}^{k-1}[0]\leftarrow\mathbf{y}_{0}^{*}$, $\mathbf{y}^{k-1}[T]\leftarrow\mathbf{y}_{f}$
      Recover controls: $\mathbf{u}_{t}^{0}=f_{\phi}(\mathbf{y}_{t}^{0},\mathbf{y}_{t+1}^{0})$
@@
Generate $\mathbf{y}_{\text{update}}^{0}=f_{\text{forward}}(\mathbf{u}_{0:T}^{0},\mathbf{y}_{0}^{*})$
Augment dataset: $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{(\mathbf{y}_{\text{update}}^{0},\mathbf{u}_{0:T}^{0})\}$
@@ Appendix B Detailed System and Dataset Description
B.1 Burgers Dynamics
@@
We follow instructions in (Wei et al., 2024) to generate a 1D Burgers’ equation dataset. Specifically, for numerical simulation, we discretized the spatial domain [0,1] and temporal domain [0,1] using the finite difference method (FDM). The spatial grid consisted of 128 points, while the temporal domain was divided into 10000 timesteps. We initiated the system with randomly sampled initial conditions and control inputs drawn from specified probability distributions. This setup allowed us to generate 90000 trajectories for training and 50 trajectories for testing purposes.
B.2 Kuramoto Dynamics
@@
B.3 Inverted Pendulum Dynamics
@@ C.1 Implementation of SEDC
  • The temporal U-Net (1D-Unet) (Janner et al., 2022) in the denoising network consists of a U-Net structure with $4$ repeated residual blocks. Each block comprises two temporal convolutions, followed by group normalization, and a final Mish nonlinearity. The channel dimensions of the downsample layers are ${1,2,4}*statedimension$. Timestep embedding is produced by a Sinusoidal Positional Encoder, following a 2-layer MLP, and the dimension of this embedding is 32. The dimension of condition embedding is the same as the system state dimension.
@@ D.3 Diffusion-based methods
    A novel approach that reformulates sequential decision-making as a conditional generative modeling problem rather than a reinforcement learning task. The core methodology involves modeling policies as return-conditional diffusion models, enabling direct learning from offline data without dynamic programming. The model can be conditioned on various factors including constraints and skills during training.
@@
Appendix E Numeric Results of Figure 2
Table 6: Performance comparison of different models across three datasets. Lower values indicate better performance for both metrics. Best and second-best results of each row are highlighted in bold and underlined respectively.
@@ Appendix E Numeric Results of Figure 2
Target Loss $J$(Energy) Target Loss $J$(Energy) Target Loss $J$(Energy)
@@
We leverage 2-D plots in the main paper to better illustrate the performance comparison of all the methods. Here we provide the provides the corresponding numerical results in detail in Table 6.
Appendix F Results and Discussion of the ablation study on GSF
Table 7: Energy consumption reduction across different dynamical systems after each round of GSF. The “Before” column shows energy performance of the trained model before finetuning. For each round, “$\Delta$%” represents the percentage reduction compared to its previous stage. Lower energy values indicate better system performance. All systems demonstrate significant initial improvements (1st round) followed by smaller incremental gains (2nd round).
System Before 1st round 2nd round
Value $\Delta$Value $\Delta$%
Burgers 8.39 5.23 37.7% 5.01 4.2%
Kuramoto 17.48 15.20 13.0% 14.90 2.0%
Swing 0.27 0.24 13.4% 0.23 2.4%
IP 9.00e-3 9.50e-4 89.4% 8.90e-4 6.3%
Average $\Delta$38.4% 3.7%
To validate GSF’s effectiveness in guiding the model toward learning the optimal (energy- efficient) target distribution, we conduct ablation studies on fine-tuning rounds and test control signal energy on the test set. The result is provided in Table 7. Before finetuning, energy performance is the poorest because of the non-optimality of the initial training samples, and the first round of GSF greatly supplemented the training toward optimality, as the energy of the produced control inputs decreases at an average of 38.4%. The increases observed in the second round persist but are significantly lower than those in the first round. This may be because the first round of GSF has already captured the most significant deviations toward optimality, while subsequent rounds primarily refine these improvements with diminishing returns. The slowing rate of improvement suggests the model is approaching a convergence point in its exploration of the optimal control space. Overall, the above result demonstrates the effectiveness of GSF in enabling exploration beyond initial training data and facilitating convergence toward optimal control strategies.
Appendix G Visualization

### v2

@@ 5.4 Ablation Study
Overall ablation study. We explore the main performance against each ablation of the original SEDC. Specifically, w/o DSD removes the inverse dynamics, unifying the diffusion of system state and control input, i.e. $\mathbf{x}=[\mathbf{u},\mathbf{y}]$. Therefore, the diffusion model is required to simultaneously capture the temporal information and implicit dynamics of the control and system trajectory. Note that the inpainting mechanism and gradient guidance are retained. w/o DMD removes the decomposition design, resulting in a single 1-D Unet structure as the denoising network, following DecisionDiff. Finally, w/o GSF reports the performance without guided self-finetuning process, which means the model only uses the original dataset to train itself. To show the sample-efficiency performance, we also investigate the results under less amount of training sample(10%). For w/o DMD and w/o DSD, we adjust the number of trainable parameters at a comparable level against the original version.
@@
Effectiveness of DSD. To evaluate DSD’s effectiveness against the curse of dimensionality, we compared the performance of original and w/o DSD models across Kuramoto systems with dimensions ranging from $N=4$ to $N=8$. Experimental results (Figure 4(a)) show that performance degradation from w/o DSD increases proportionally with system dimensionality, demonstrating DSD’s enhanced effectiveness in higher-dimensional systems and validating its capability to address dimensionality challenges. Further, we can observe that the decrease changes more rapidly with increasing dimensionality compared to Kuramoto, indicating that DSD exhibits heightened sensitivity to dimensionality in higher-dimensional systems. We also investigate the effectiveness of dynamical learning in Appendix G.4.
Table 2: Performance degradation using different denoiser output with varying nonlinearity strength $\gamma$ in the Kuramoto system.
@@
Effectiveness of DMD. To investigate the contribution of DMD’s dual-Unet architecture to nonlinearity learning, we conducted experiments on the Kuramoto system with varying degrees of nonlinearity (controlled by the coefficient $\gamma\in\{1,2,4\}$ of the nonlinear sinusoidal term, where larger values indicate stronger nonlinearity). We compared the performance between using only the linear intermediate output ($\hat{\mathbf{x}}_{0}=\mathbf{O}_{1}$) of the denoising network and the original nonlinear output ($\hat{\mathbf{x}}_{0}=\mathbf{O}_{1}+\mathbf{O}_{2}$) in terms of Target Loss. The Dec. indicates the reduction in target loss achieved by nonlinear output $\mathbf{O}_{1}+\mathbf{O}_{2}$ compared to linear output $\mathbf{O}_{1}$. As shown in Table 2, the magnitude of loss reduction increases proportionally with the nonlinearity strength $\gamma$, indicating that the quadratic term exhibits enhanced capability in capturing nonlinear dynamics as the system’s nonlinearity intensifies. This comparison confirms that the nonlinear branch $\mathbf{O}_{2}$ is essential for capturing dynamics as nonlinearity ($\gamma$) increases. The superiority of the decomposition architecture itself is further evidenced in Table 1, where w/o DMD shows consistently higher target loss. We also show that DMD is sufficient for modeling even higher-order dynamics in Appendix G.5.
Effectiveness of GSF. To validate GSF’s ability to refine the control law towards lower energy cost, we tracked test-set energy consumption across finetuning rounds. As shown in Figure 4(b), the first GSF round yields a dramatic improvement over the initial model, reducing energy by 37.7% (Burgers) and 13.0% (Kuramoto). Subsequent rounds offer minor refinements (4.2% and 2.0% respectively), indicating convergence to a near-optimal law. This confirms GSF effectively guides the model beyond its initial suboptimal training data to discover more energy-efficient control solutions.
Figure 5: Comparison on a rank-deficient system showing SEDC’s robustness to multi-modal control.
Figure 6: Stabilization performance on the 39-bus New England Grid comparing SEDC with diffusion baselines.
Robustness to Non-invertibility. A critical question is whether DSD fails when multiple controls map to the same state transition (multi-modality). To investigate this, we tested SEDC against the joint-diffusion baseline (w/o DSD) on a rank-deficient system where control is non-unique. Details can be found in Appendix G.3. As shown in Figure 6, SEDC outperforms the baseline by reducing Target Loss by over 70%. We attribute this success to task decomposition: joint baselines struggle to model the complex multi-modal joint distribution $p(\mathbf{y},\mathbf{u})$, often leading to mode averaging. In contrast, SEDC first generates a coherent state trajectory $\mathbf{y}_{0:T}$. Conditioned on this valid path, the inverse model $f_{\phi}$ transforms the intractable joint generation problem into a simplified supervised task: identifying a feasible control (e.g., the smoothest one) that satisfies the transition.
5.5 Case Study
Real-world Power Grid. To validate scalability on real-world infrastructure, we applied SEDC to the 39-bus New England power grid (Swing Equation) (Baggio et al., 2021), a real-world benchmark for post-fault stabilization (setup details in Appendix F). As shown in Figure 6, SEDC achieves a 60.5% reduction in Target Loss compared to the strongest baseline (DecisionDiffuser) while maintaining lower energy cost. This significant margin demonstrates SEDC’s superior ability to capture the highly coupled nonlinearities of power networks, where our global planning approach avoids the compounding errors typical of myopic methods in recovering synchronous states.
@@ 6 Conclusion
In this paper, we presented SEDC, a novel sample-efficient diffusion-based framework for complex nonlinear system control. By synergistically integrating Decoupled State Diffusion (DSD), Dual-Mode Decomposition (DMD), and Guided Self-finetuning (GSF), SEDC achieves superior control performance with remarkable data efficiency. Our experiments show that SEDC can match state-of-the-art accuracy using just 10% of the training data. The framework’s robustness is further validated through rigorous testing on high-dimensional PDEs and real-world systems, confirming the broad applicability and effectiveness of our design principles. These results mark a significant advancement in developing practical and sample-efficient solutions for complex system control.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Algorithm form of SEDC
Predict clean states: $\hat{\mathbf{y}}^{k}=\mathcal{G}_{\theta}(\mathbf{y}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})$
Predict controls: $\tilde{\mathbf{u}}_{t}=f_{\phi}({\mathbf{y}}_{t},{\mathbf{y}}_{t+1})$
Compute losses:  $L_{\text{diff}}=\|\mathbf{y}-\hat{\mathbf{y}}^{k}\|^{2}$
@@
      Predict $\hat{\mathbf{y}}^{k}=\mathcal{G}_{\theta}(\mathbf{y}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})$
Compute gradient: $g=\nabla_{\mathbf{y}^{k}}J(\tilde{\mathbf{y}}^{k},\hat{\mathbf{u}}^{k})$, where $\tilde{\mathbf{u}}_{t}^{k}=f_{\phi}(\hat{\mathbf{y}}_{t}^{k},\hat{\mathbf{y}}_{t+1}^{k})$
@@
Enforce constraints: $\mathbf{y}^{k-1}_{0}\leftarrow\mathbf{y}_{0}^{*}$, $\mathbf{y}^{k-1}_{T}\leftarrow\mathbf{y}_{f}$
  Recover controls: $\tilde{\mathbf{u}}_{t}^{0}=f_{\phi}(\mathbf{y}_{t}^{0},\mathbf{y}_{t+1}^{0})$
@@
Generate $\tilde{\mathbf{y}}_{[0:T]}^{0}=f_{\text{forward}}(\tilde{\mathbf{u}}_{[0:T]}^{0},\mathbf{y}_{0}^{*})$
Augment dataset: $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{(\tilde{\mathbf{y}}_{[0:T]}^{0},\tilde{\mathbf{u}}_{[0:T]}^{0}\}$
@@ Appendix B Detailed System and Dataset Description
B.1 Overall Introduction
Our benchmark selection (Inverted Pendulum, Kuramoto, and Burgers) follows established control systems research practice, chosen for real-world relevance and diverse nonlinearity and complexity:
  • Inverted Pendulum: state 2, control 1, timestep 128
  • Kuramoto: state 8, control 8, timestep 15
  • Burgers: state 128, control 128, timestep 10
B.2 Burgers Dynamics
@@
We follow instructions in Wei et al. (2024b) to generate a 1D Burgers’ equation dataset. Specifically, for numerical simulation, we discretized the spatial domain [0,1] and temporal domain [0,1] using the finite difference method (FDM). The spatial grid consisted of 128 points, while the temporal domain was divided into 10000 timesteps. We initiated the system with randomly sampled initial conditions and control inputs drawn from specified probability distributions. This setup allowed us to generate 90000 trajectories for training and 50 trajectories for testing purposes.
B.3 Kuramoto Dynamics
@@
B.4 Inverted Pendulum Dynamics
@@ C.1 Implementation of SEDC
  • The temporal U-Net (1D-Unet) (Janner et al., 2022) in the denoising network consists of a U-Net structure with $4$ repeated residual blocks. Each block comprises two temporal convolutions, followed by group normalization, and a final Mish nonlinearity. The channel dimensions of the downsample layers are 1,2,4statedimension{1,2,4}*statedimension. Timestep embedding is produced by a Sinusoidal Positional Encoder, following a 2-layer MLP, and the dimension of this embedding is 32. The dimension of condition embedding is the same as the system state dimension.
@@ D.3 Diffusion-based methods
    A novel approach that reformulates sequential decision-making as a conditional generative modeling problem rather than a reinforcement learning task. The core methodology involves modeling policies as return-conditional diffusion models, enabling direct learning from data without dynamic programming. The model can be conditioned on various factors including constraints and skills during training.
@@
Appendix E Detailed Results of Figure 2
Table 5: Performance comparison of different models across three datasets. TL (Target Loss) and $J$ (Energy) are reported, with lower values indicating better performance for both metrics. We report the mean and the standard error over 5 random seeds. Following previous work(e.g. Ajay et al. (2022)), we highlight the best-performed results in bold.
@@ Appendix E Detailed Results of Figure 2
TL $J$ TL $J$ TL $J$
@@
We leverage 2-D plots in the main paper to better illustrate the performance comparison of all the methods. Here we provide the provides the corresponding numerical results in detail in Table 5 and pareto frontiers in Fig 7. Results confirm our method maintains competitive performance across all datasets.
Figure 7: Pareto frontier of target loss and energy cost $J$ across different datasets of our method and two SOTA baselines DecisionDiffuser and DiffPhyCon. The closer the data point is to the bottom left, the better the performance.
Figure 8: Illustration of the power grid swing dynamics control task. The system simulates a recovery scenario following a line fault, where the objective is to compute an optimal control sequence that drives the generators from a disturbed, un-synchronized state back to steady-state operation (zero deviation). Several component of this figure are adapted from Baggio et al. (2021).
Appendix F Power Grid Swing Dynamics Details
In Section 5.5, we presented comparative results on the power grid swing dynamics. Here we provide the specific experimental configuration.
System Description. The Swing Equation system models the rotational dynamics of synchronous generators in a power network. We utilize the New England 10-generator 39-bus system. The state space consists of the rotor angles and frequencies of the generators, resulting in a high-dimensional nonlinear system.
Task Setup. Following the experimental design in Baggio et al. (2021), the control task simulates a recovery scenario following a line fault. The process is illustrated in Figure 8. The objective is to compute an optimal control sequence that drives the generators from a disturbed, un-synchronized state back to the steady-state operation (zero deviation). The trajectories are sampled with 18 state variables and 9 control inputs over a horizon of 32 timesteps. This setup incorporates realistic physical parameters and disturbances, serving as a rigorous test for controlling coupled nonlinear dynamics.
Appendix G Supplementary Experiments
G.1 Comparison against MPC
Table 6: Comparison of SEDC (Ours) and Model Predictive Control (MPC) across three control tasks. Results show target loss (MSE), control cost (J), and inference time in seconds. We implemented the MPC that uses a neural network architecture with residual connections to learn system dynamics from data, then solves finite-horizon optimization problems at each timestep using the gradient of the summation of target loss and control energy, which iteratively refines the sampling distribution toward optimal control sequences. The control horizons of Kuramoto,Burgers and IP are 15, 10 and 128.
Model Kuramoto Burgers IP
Target loss J Time(s) Target loss J Time(s) Target loss J Time(s)
MPC 9.10e-03 0.34 $\sim$50 1.89e-01 0.001 $\sim$30 4.33e-04 1.01e-03 >>1000
Ours 8.90e-06 14.9 $\sim$1.5 9.80e-05 5.01 $\sim$2.5 3.49e-04 8.90e-04 $\sim$0.5
We additionally compare SEDC with data-driven Model Predictive Control (MPC). We implement MPC in a data-driven way, where we train an MLP for the forward model. Results in Table 6 show MPC achieves higher target losses across all tasks, likely due to error accumulation. MPC also has significantly longer inference times (e.g. ¿1000 vs. 0.5) that increase with control horizon. SEDC directly maps initial/target states to complete control trajectories, avoiding compounding errors and reducing computation time.
G.2 Validation on High-Dimensional 2D PDE Control: Jellyfish Locomotion
To address the critical question of our method’s scalability and effectiveness on truly high-dimensional control problems, we conducted additional experiments on a challenging 2D PDE control benchmark: the locomotion of a jellyfish. This task represents a state-of-the-art challenge in data-driven control of physical systems, involving complex fluid-solid interactions governed by the 2D incompressible Navier-Stokes equations (Wei et al., 2024b). Unlike the systems in the main text, this benchmark provides a testbed with a significantly higher state-space dimension, allowing for a rigorous evaluation of SEDC’s scalability.
Experimental Setup The objective is to control the opening angle of the jellyfish’s wings to achieve a target locomotion pattern. The system state is a high-dimensional PDE field representing the fluid velocity and pressure, while the control input is a scalar time series representing the wing angle.
  • State Representation: Each state at a given timestep is represented by a $3\times 32\times 32$ tensor, resulting in a state dimension of 3,072. This is a substantial increase in complexity compared to the 1D systems.
  • Dataset: We generated a dataset of 20,000 control trajectories for training, following the standard procedure for this benchmark.
Overall Performance Comparison We first evaluated SEDC against strong diffusion-based baselines using the full training dataset. The results, shown in Table 7, assess control accuracy (Target Loss), energy efficiency (Energy), and computational cost (Training and Inference Time).
Table 7: Performance comparison on the 2D Jellyfish Locomotion control task (100% training data). SEDC achieves the best control accuracy with a competitive computational profile.
Model Target Loss ($\downarrow$) Energy ($\downarrow$) Train Time (hrs) Inference Time (s)
DecisionDiffuser 1.74e-4 $\pm$ 7.4e-5 2.016 $\pm$ 0.08 3.0 6.0
AdaptDiffuser 1.77e-4 $\pm$ 1.7e-5 2.001 $\pm$ 0.71 3.0 6.0
DiffPhyCon 1.77e-4 $\pm$ 5.4e-5 2.157 $\pm$ 0.21 4.0 8.0
SEDC (Ours) 1.70e-4 $\pm$ 1.3e-5 2.015 $\pm$ 0.43 3.0 6.5
The results demonstrate that SEDC achieves the best control accuracy (lowest Target Loss) among all methods, confirming that its architectural advantages for sample-efficient learning translate effectively to high-dimensional PDE systems. Furthermore, its computational cost remains on par with the fastest baselines, highlighting its practicality.
Sample Efficiency Analysis A key claim of our work is superior sample efficiency. To specifically validate this on a high-dimensional task, we compared the performance of SEDC against the strongest baseline (DecisionDiffuser) when trained on only 20% of the available data versus the full dataset.
Table 8: Sample efficiency comparison on the Jellyfish task. Results show the final Target Loss. SEDC trained on only 20% of the data outperforms the baseline trained on 100% of the data.
Model 100% Training Data 20% Training Data
DecisionDiffuser 1.74e-4 4.52e-4
SEDC (Ours) 1.70e-4 1.75e-4
As shown in Table 8, SEDC exhibits remarkable sample efficiency. When trained on just 20% of the data, its performance is statistically on par with the baseline trained on the full dataset (1.75e-4 vs 1.74e-4). In contrast, the baseline’s performance degrades significantly when data is limited. This comprehensive validation on a high-dimensional PDE benchmark strongly supports our paper’s central claim: SEDC provides a scalable and highly sample-efficient framework for the control of complex nonlinear systems.
G.3 System Applicability and Non-Invertible Dynamics Details
In Section 5.5, we demonstrated SEDC’s robustness on non-invertible systems. Here we provide the specific dynamics configuration.
Rank-Deficient Linear System Setup. We constructed a linear system with a rank-deficient control matrix to simulate control redundancy (infinite control solutions):
$\dot{\mathbf{y}}=\mathbf{A}\mathbf{y}+\mathbf{B}\mathbf{u},\quad\mathbf{A}=\begin{bmatrix}0&1\\ -1&-0.5\end{bmatrix},\quad\mathbf{B}=\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ (10)
where $\mathbf{y}\in\mathbb{R}^{2}$ and $\mathbf{u}\in\mathbb{R}^{2}$. The matrix $\mathbf{B}$ has rank 1, meaning the controls $u_{1}$ and $u_{2}$ are coupled ($u_{1}+u_{2}$ affects the system), making individual control recovery mathematically impossible. Despite this, as reported in the main text, SEDC successfully learns a valid control law (implicitly learning an equal distribution or minimum norm solution), whereas the joint diffusion baseline fails to converge to a consistent state-control pair.
Additional Experiment: Non-Affine MIMO System. We also verified performance on a nonlinear non-affine system: $\dot{x}_{1}=-x_{1}+x_{2}$, $\dot{x}_{2}=\sin(x_{1})-0.5x_{2}+u_{1}^{2}-u_{2}^{2}$. SEDC similarly achieved a Target Loss of 5.0e-2 compared to 8.1e-2 for the baseline, further confirming robustness.
G.4 The effectiveness of dynamical learning
To investigate the effectiveness of dynamical learning, we compared the consistency between action sequences and diffusion-sampled state trajectories in models with and without DSD. While both approaches can sample state trajectories from diffusion samples, they differ in action generation: SEDC uses inverse dynamics prediction, whereas w/o DSD obtains control signals directly from diffusion samples by simultaneously diffusing states and control inputs. We test both models using identical start-target conditions and visualize the state induced from the generated inputs and the state sampled from the diffusion model, along with the difference (error) between the above two states in Figure 9. We can observe that SEDC’s action-induced state trajectories showed significantly higher consistency with sampled trajectories compared to w/o DSD, demonstrating that DSD using inverse dynamics achieves more accurate learning of control-state dynamical relationships.
G.5 Sufficiency of the 2nd-Order DMD Architecture
To address whether our 2nd-order DMD approximation is sufficient for systems with richer nonlinearities, we conducted a targeted experiment on synthetic 2D systems with controlled nonlinear terms. We designed three systems: one purely linear, one with a quadratic term, and one with a cubic term. We compared our full DMD model against a single-UNet baseline and a linear-only ablation of our model.
Table 9: Performance (Target Loss) on synthetic systems with varying orders of nonlinearity. Our 2nd-order DMD model is sufficient to effectively control the 3rd-order system.
System Dynamics Single-UNet DMD (Ours) Linear-Only
1st-Order (Linear) 1.05e-6 8.68e-7 1.54e-6
2nd-Order (Quadratic) 6.63e-6 5.20e-6 2.50e-4
3rd-Order (Cubic) 7.00e-5 6.43e-5 5.80e-3
As shown in Table 9, our DMD model achieves the lowest error on the 3rd-order system, demonstrating its sufficiency. The performance of the Linear-Only model is nearly 90x worse, confirming that the nonlinear branch ($\mathbf{O}_{2}$) is not redundant and is critical for capturing the system’s dynamics. This principled, sample-efficient design proves robust even for complex systems beyond its explicit Taylor-expansion motivation.
G.6 Robustness to Observation Noise
To evaluate the robustness of our method, a critical factor for real-world applicability, we conducted new experiments to validate SEDC’s performance when trained on data corrupted by observation noise. This setup simulates practical scenarios where state measurements are imperfect.
Experimental Setup We added zero-mean Gaussian noise with varying standard deviations ($\sigma$) to the state observations in the training data for both the Kuramoto and Burgers systems. We then retrained our model and the strongest baselines from scratch on this noisy data and evaluated their control accuracy on a clean, noise-free test set.
Results The results, summarized in Table 10, show that SEDC consistently achieves superior control accuracy across all noise levels.
Table 10: Performance (Target Loss) comparison on noisy training data. Energy cost is shown in parentheses. SEDC demonstrates consistently higher accuracy under noisy conditions.
System Noise ($\sigma$) SEDC (Ours) AdaptDiffuser DiffPhyCon
Kuramoto 0 8.90e-6 (14.90) 1.76e-5 (26.23) 4.80e-4 (18.72)
0.001 1.21e-5 (17.19) 8.00e-4 (16.24) 3.95e-3 (11.50)
0.01 5.61e-4 (10.40) 1.75e-3 (3.37) 3.24e-3 (17.58)
0.1 3.14e-3 (0.42) 3.96e-3 (0.98) 3.14e-3 (2.22)
Burgers 0 9.80e-5 (5.01) 2.28e-4 (4.65) 1.62e-4 (5.15)
0.001 1.86e-4 (5.01) 2.81e-4 (4.73) 6.05e-4 (5.11)
0.01 2.25e-4 (4.40) 3.98e-4 (4.21) 8.69e-4 (4.58)
0.1 6.97e-4 (4.25) 1.22e-3 (3.75) 2.89e-3 (3.13)
Notably, on the Burgers system with medium noise ($\sigma=0.01$), SEDC’s target loss of 2.25e-4 is 1.8x better than AdaptDiffuser’s and 3.9x better than DiffPhyCon’s. This superior robustness stems from our key architectural innovations. DMD’s decomposition helps capture the core system dynamics resiliently, while DSD’s focus on learning a smoother state-only distribution prevents overfitting to noise, a common issue when modeling complex joint state-control distributions.
Appendix H Visualization